# Chiral orbital current driven topological Hall effect in Mn3Si2Te6

### v2

@@ Abstract
Introduction:— The interplay of topology, magnetism, and orbital degrees of freedom has emerged as a central theme in condensed matter physics, giving rise to a variety of unconventional transport phenomena. In particular, the topological Hall effect (THE) is widely regarded as a transport fingerprint of nontrivial spin chirality and Berry curvature associated with real-space spin textures. With the discovery of the skyrmion lattice in the chiral ferromagnet MnSi Muhlbauer2009, THE has emerged as a platform to explore exotic physical properties. In recent times, THE has been observed in many non-centrosymmetric systems Neubauer2009; Kanazawa2011; Surgers2014; Gallagher2017, and in a few centrosymmetric systems as well Wang2016; Zheng2021; Rout2019; Kurumaji2019. In most systems, THE is shown to originate from broken inversion symmetry giving rise to the Dzyaloshinsky-Moriya interaction (DMI) which favours the formation of topologically nontrivial spin textures like skyrmions, anti-skyrmions, skyrmonium, (anti)merons, bimerons, etc Nayak2017; Zhang2018; Gao2020; Yu2018; Nagase2021; Muhlbauer2009. Recent theoretical developments, however, suggest that orbital degrees of freedom alone can generate topological–chiral interactions through emergent orbital magnetism associated with chiral orbital currents (COC), independent of conventional spin–orbit coupling (SOC) PhysRevB.98.195439; Grytsiuk2020. Such orbital-driven interactions have been predicted to stabilize exotic magnetic textures and to engender new routes toward topological transport responses Grytsiuk2020; Yang2023MonopoleOrbital; Go2018Intrinsic; Nikolaev2024ChiralOrbital; dosSantosDias2016ChiralityDrivenOrbital; PhysRevResearch.5.043052.
@@
COC has been under investigation since the early 2000s with studies on high-TC cuprates Varma2006; Varma2014; Bourges2021; DiMatteo2012; Pershoguba2013; Scagnoli2011; Pershoguba2014; Yakovenko2015, iridates Zhao2015; Jeong2017; Murayama2021, Kagome superconductors and magnets Jiang2021; Feng2021; Teng2022 and moiré heterostructures Serlin2020; Tschirhart2021; Sharpe2019. Later it was also shown that the coupling of COC with long-range magnetic order can break both mirror and time-reversal symmetry and can lead to interesting phenomena like colossal magnetoresistance (CMR) Zhang2022ChiralOrbitalCurrents. Coupling the COC with applied current can also present intriguing physics, especially giving rise to a current sensitive unique Hall effect in Mn3Si2Te6 (MST) Zhang2024CurrentSensitiveHall. In MST, the COC flows along the Te edges in the ab-plane, giving rise to an orbital moment and a real-space emergent magnetic field Bc, both directed along the c-axis. Currently, most of the studies on MST are directed towards the unique phenomena of CMR Seo2021ColossalAMR; PhysRevB.111.174419; PhysRevB.106.045106; PhysRevB.103.L161105; Zhang2024; Lovesey2023 with very few findings on the Hall resistivity of the same. The observation of the unique the Hall effect in MST cannot be completely explained away by considering the contributions of ordinary Hall effect (OHE) and anomalous Hall effect (AHE) Zhang2024CurrentSensitiveHall; PhysRevB.108.125103. Thus, the complete picture of the exact contributions to the Hall effect of MST and the physics behind it is still unknown.
@@
Experimental Methods:— Single crystals of Mn3Si2Te6 (MST) have been prepared using the chemical vapour transport (CVT) method, where iodine (I2) is used as the transport agent (see Sec. I of SM SM for details on sample preparation). The typical crystals were of 1-2 mm in size. Energy dispersive spectroscopy (EDS) measurements were performed to confirm the elemental composition of the grown crystals. X-ray diffraction (XRD) measurements were performed using a PANalytical Empyrean diffractometer. Both powder XRD and single crystal XRD measurements were carried out to confirm the purity and crystallinity of the grown crystals. The powder XRD data was profile fitted using Pm space group symmetry using FullProfSuite software. The temperature dependence of magnetization and the isothermal magnetization hysteresis measurements of the bulk single crystals have been carried out using a Quantum Design physical properties measurement system (PPMS) using the standard vibrating sample magnetometry (VSM) technique. For magento-transport measurements, we use a variable temperature insert (VTI) cryostat provided by CRYOGENIC Inc., UK. The raw MR and Hall curves were symmetrized and antisymmetrized, respectively, to eliminate the effects of electrode misalignment.
@@
Results and Discussion:— The single crystal XRD data of the bulk single crystal reveals the (001) orientation of the crystal surface (refer to Sec. I of SM SM for the details). The powder XRD spectra (see Fig. 1(a) of PhysRevB.111.174419 for details) reveals that the grown single crystals are of high purity. Additionally, the EDS spectra (refer to Sec. I of SM SM for the details) shows prominent peaks of Mn, Si and Te, confirming that the obtained stoichiometry is close to the targeted one.
@@
Figure 1(a) shows the temperature-dependent magnetization M(T) of bulk MST single crystals at an applied field of 0.1 T. It shows that the ab-plane is the easy plane where the magnetization saturates at around 1.5 /Mn, while the c-axis is the hard axis. The Curie temperature as obtained from Fig. 1(a) is close to 78 K, which is consistent with earlier reports PhysRevB.111.174419; PhysRevB.106.045106; PhysRevB.103.245122; PhysRevB.108.125103; PhysRevB.103.L161105; Zhang2024CurrentSensitiveHall; Zhang2022ChiralOrbitalCurrents; Seo2021ColossalAMR; PhysRevB.95.174440. The inset of Fig. 1(a) shows the isothermal magnetization at 5 K for both the crystallographic axes, suggesting that the c-axis magnetization does not saturate even at a higher field of 5 T. The temperature dependence of electrical resistivity for the bulk single crystals of MST is shown in Fig. 1(b), which indicates its semiconducting nature similar to other reports PhysRevB.106.045106; PhysRevB.103.245122; PhysRevB.108.125103. The inset of Fig. 1(b) shows the field dependence of magnetoresistance (MR) for bulk MST samples at 5 K with B c showing the well-known CMR effect.
@@
Figure 1(c) shows the field dependence of Hall resistivity () at 5 K for the bulk MST single crystal. The appearance of a distinct sharp peak in the low field region of the Hall resistivity data suggests additional contributions to the data apart from the OHE and AHE, similar to what was observed in previous reports PhysRevB.108.125103; PhysRevB.103.L161105; Zhang2024CurrentSensitiveHall. This indicates possible contributions from the THE. Similar peaks in the field-dependent Hall resistivity of the nanoflake MST device are observed in Fig. 1(d). For the device, the peak becomes sharper and more prominent, indicating that the THE contribution possibly enhances with decreasing thickness.
@@
To further examine the origin of the sharp peak observed in the Hall resistivity data, we measure the field dependence of at different temperatures for both bulk and nanoflake device of MST, as shown in Fig. 2(a) and 2(b), respectively. For bulk, the peak in the Hall resistivity data persists up to to 70 K, while for the device, it starts to die out above 30 K. The occurrence of this peak in the curve for bulk MST has been attributed to COC in previous reports Zhang2024CurrentSensitiveHall. This COC is known to induce an orbital moment (MCOC) along the c axis Zhang2022ChiralOrbitalCurrents, that underpins the well-known CMR in bulk MST. This MCOC produces a magnetic field Bc in the real space, which is directed along the c-axis. When B c, it adds up to the external field applied and influences the Hall resistivity Zhang2024CurrentSensitiveHall.
@@
where , and , refers to ordinary, anomalous and topological Hall contribution to , respectively. Ms refers to saturation magnetization, while R0 and are the ordinary and anomalous Hall coefficients, respectively. The ordinary Hall component usually dominates as a high-field linear component of the data and is separated out by calculating R0 using a linear fit at high fields. The anomalous Hall component is isolated by fitting with a step-like function. Figure 2(c) and (d) shows the fitted curves for bulk and device MST, respectively. The red solid line shows the ordinary and anomalous contributions in with B c at 5 K in both cases. The sharp peak region, in excess of the ordinary and anomalous Hall contribution, is attributed to the THE and is extracted by subtracting the curves by the red solid line data.
@@
The contour plots of the extracted THE amplitude for both the bulk and nanoflake device are shown in Fig. 2(e),(f), respectively. From the figures, it is evident that the THE amplitude for the bulk persists up to a higher temperature when compared to the nanoflake device. This could be because of the COC state being more vulnerable to thermal fluctuations when one approaches nanoscale thickness limit compared to the bulk. The THE amplitude mimicking the sensitivity of the COC state to sample thickness suggests that the former might be governed by the latter.
@@
To ascertain the causal connection of the COC and the observed THE in MST, we perform the magnetic field dependence of the Hall resistivity as a function of the applied current. Since it is well known that the strength of the COC decreases with increasing current Zhang2024CurrentSensitiveHall, the relation of to the current will be crucial in studying its actual origin. Figure 3 depicts the extracted THE amplitudes as a function of temperature and magnetic field for different applied currents for both bulk [Fig 3(a) - (e)] and device [Fig 3(f) - (j)], respectively. It is evident from the figure that upon increasing the current, the THE amplitudes decrease for both the bulk and the nanoflake device.
@@
For bulk MST, the THE amplitude remains non-zero even up to 70 K when the applied current is small (down to 50 A) as can be seen from Fig. 3(a). As the current increases to 500 A, the region of the non-zero THE amplitude splits into islands of topological Hall signals [Fig. 3(b)], which shrink and eventually vanish upon increasing the current to 1 mA. The THE amplitude persists only up to 30 K for 1 mA, and keeps decreasing down to lower temperatures with increasing current. Finally, for 10 mA, the THE signal disappears completely as seen in Fig. 3(e), similar to that reported in Zhang2024CurrentSensitiveHall. This is expected as the COC is stronger for lower currents but on increasing the current it starts to weaken. At high current, the COC state () is destroyed completely, making way for a trivial state (). This similar trend is seen to be followed by the THE signal of the bulk MST crystals, indicating that the two might be interconnected.
Although the THE signal from the nanoflake device of MST replicates the exact current dependence as the bulk MST, the device shows the THE behaviour at a much lower current than the bulk as is evident from, Fig. 3(f) - (j). At lower thickness, the cross sectional area becomes much smaller which leads to a much higher current density for a given current (). Thus, for a similar current, the orbital motions leading to COC is expected to be suppressed more for thin flakes than for the bulk. Hence, the THE amplitude for the nanoflake device dies out at a much lower current than that of the bulk due to much higher current density.
@@
From previous reports, it is well-established that the CMR portrayed by the MST is governed primarily by the COC Zhang2022ChiralOrbitalCurrents; Zhang2024CurrentSensitiveHall; PhysRevB.111.174419. Thus, drawing a correlation between the THE and the magnetoresistance of MST in bulk and device, respectively, will provide a firm assurance of the COC origin of the THE. Figure 4(a),(b) shows the temperature dependence of MR and relative topological Hall conductivity for the bulk and device of MST, respectively. The MR is calculated as, , where, with being the magnetoresistivity measured at an applied magnetic field of B T, and the zero-field (ZF) resistivity. The hall conductivity is given by , where and are the longitudinal and Hall resistivity, respectively. Similarly, the topological Hall conductivity is given by , where is the extracted topological Hall resistivity.
@@
Thus, we observe that the THE in both bulk and nanoscale device of MST has a very strong correlation with the phenomena of CMR. As the latter is directly influenced by the COC in MST, the former could possibly have the same origin. It has already been reported that the COC develop a moment MCOC which induces a magnetic field Bc in real space, directed along the c-axis. This gives rise to a non-zero local scalar spin chirality, and the moment MCOC acts as the topological orbital moment (TOM). This couples with the local spins of Mn giving rise to spin-chiral interaction (SCI), mediated by SOC. Also, each individual COC can interact with each other, giving rise to a chiral-chiral interaction (CCI) Grytsiuk2020. The formation of these two topological chiral interactions give rise to a finite spin chirality. Now, the THE contribution to the resistivity comes from two terms: , where refers to the transverse charge current due to chiral orbital textures and comes from chiral spin textures PhysRevB.98.195439. Hence, systems with strong dominating chiral orbital textures, like COC, can also show THE, which might be the reason behind the observed THE in MST.

### v3

@@ Abstract
Introduction:— The interplay of topology, magnetism, and orbital degrees of freedom has emerged as a central theme in condensed matter physics, giving rise to a variety of unconventional transport phenomena. In particular, the topological Hall effect (THE) is widely regarded as a transport fingerprint of nontrivial spin chirality and Berry curvature associated with real-space spin textures. With the discovery of the skyrmion lattice in the chiral ferromagnet MnSi Mühlbauer et al. (2009), THE has emerged as a platform to explore exotic physical properties. In recent times, THE has been observed in many non-centrosymmetric systems Neubauer et al. (2009); Kanazawa et al. (2011); Sürgers et al. (2014); Gallagher et al. (2017), and in a few centrosymmetric systems as well Wang et al. (2016); Zheng et al. (2021); Rout et al. (2019); Kurumaji et al. (2019). In most systems, THE is shown to originate from broken inversion symmetry giving rise to the Dzyaloshinsky-Moriya interaction (DMI) which favours the formation of topologically nontrivial spin textures like skyrmions, anti-skyrmions, skyrmonium, (anti)merons, bimerons, etc Nayak et al. (2017); Zhang et al. (2018); Gao et al. (2020); Yu et al. (2018); Nagase et al. (2021); Mühlbauer et al. (2009). Recent theoretical developments, however, suggest that orbital degrees of freedom alone can generate topological–chiral interactions through emergent orbital magnetism associated with chiral orbital currents (COC), independent of conventional spin–orbit coupling (SOC) Denisov et al. (2018); Grytsiuk et al. (2020). Such orbital-driven interactions have been predicted to stabilize exotic magnetic textures and to engender new routes toward topological transport responses Grytsiuk et al. (2020); Yang et al. (2023); Go et al. (2018); Nikolaev et al. (2024); dos Santos Dias et al. (2016); Busch et al. (2023).
@@
COC has been under investigation since the early 2000s with studies on high-TC cuprates Varma (2006, 2014); Bourges et al. (2021); Di Matteo and Norman (2012); Pershoguba et al. (2013); Scagnoli et al. (2011); Pershoguba et al. (2014); Yakovenko (2015), iridates Zhao et al. (2015); Jeong et al. (2017); Murayama and other collaborators (2021), Kagome superconductors and magnets Jiang et al. (2021); Feng et al. (2021); Teng et al. (2022) and moiré heterostructures Serlin et al. (2020); Tschirhart et al. (2021); Sharpe et al. (2019). Later it was also shown that the coupling of COC with long-range magnetic order can break both mirror and time-reversal symmetry and can lead to interesting phenomena like colossal magnetoresistance (CMR) Zhang et al. (2022). Coupling the COC with applied current can also present intriguing physics, especially giving rise to a current sensitive unique Hall effect in Mn3Si2Te6 (MST) Zhang et al. (2024a). In MST, the COC flows along the Te edges in the ab-plane, giving rise to an orbital moment and a real-space emergent magnetic field Bc, both directed along the c-axis. Currently, most of the studies on MST are directed towards the unique phenomena of CMR Seo et al. (2021); Das and Mukhopadhyay (2025); Wang et al. (2022); Ni et al. (2021); Zhang et al. (2024b); Lovesey (2023) with very few findings on the Hall resistivity of the same. The observation of the unique the Hall effect in MST cannot be completely explained away by considering the contributions of ordinary Hall effect (OHE) and anomalous Hall effect (AHE) Zhang et al. (2024a); Ran et al. (2023). Thus, the complete picture of the exact contributions to the Hall effect of MST and the physics behind it is still unknown.
@@
Experimental Methods:— Single crystals of Mn3Si2Te6 (MST) have been prepared using the chemical vapour transport (CVT) method, where iodine (I2) is used as the transport agent (see Sec. I of SM 32 for details on sample preparation). The typical crystals were of 1-2 mm in size. Energy dispersive spectroscopy (EDS) measurements were performed to confirm the elemental composition of the grown crystals. X-ray diffraction (XRD) measurements were performed using a PANalytical Empyrean diffractometer. Both powder XRD and single crystal XRD measurements were carried out to confirm the purity and crystallinity of the grown crystals. The powder XRD data was profile fitted using Pm space group symmetry using FullProfSuite software. The temperature dependence of magnetization and the isothermal magnetization hysteresis measurements of the bulk single crystals have been carried out using a Quantum Design physical properties measurement system (PPMS) using the standard vibrating sample magnetometry (VSM) technique. For magento-transport measurements, we use a variable temperature insert (VTI) cryostat provided by CRYOGENIC Inc., UK. The raw MR and Hall curves were symmetrized and antisymmetrized, respectively, to eliminate the effects of electrode misalignment.
@@
Results and Discussion:— The single crystal XRD data of the bulk single crystal reveals the (001) orientation of the crystal surface (refer to Sec. I of SM 32 for the details). The powder XRD spectra (see Fig. 1(a) of Das and Mukhopadhyay (2025) for details) reveals that the grown single crystals are of high purity. Additionally, the EDS spectra (refer to Sec. I of SM 32 for the details) shows prominent peaks of Mn, Si and Te, confirming that the obtained stoichiometry is close to the targeted one.
@@
Figure 1(a) shows the temperature-dependent magnetization M(T) of bulk MST single crystals at an applied field of 0.1 T. It shows that the ab-plane is the easy plane where the magnetization saturates at around 1.5 /Mn, while the c-axis is the hard axis. The Curie temperature as obtained from Fig. 1(a) is close to 78 K, which is consistent with earlier reports Das and Mukhopadhyay (2025); Wang et al. (2022); Liu et al. (2021); Ran et al. (2023); Ni et al. (2021); Zhang et al. (2024a, 2022); Seo et al. (2021); May et al. (2017). The inset of Fig. 1(a) shows the isothermal magnetization at 5 K for both the crystallographic axes, suggesting that the c-axis magnetization does not saturate even at a higher field of 5 T. The temperature dependence of electrical resistivity for the bulk single crystals of MST is shown in Fig. 1(b), which indicates its semiconducting nature similar to other reports Wang et al. (2022); Liu et al. (2021); Ran et al. (2023). The inset of Fig. 1(b) shows the field dependence of magnetoresistance (MR) for bulk MST samples at 5 K with B c showing the well-known CMR effect.
@@
Figure 1(c) shows the field dependence of Hall resistivity () at 5 K for the bulk MST single crystal. The appearance of a distinct sharp peak in the low field region of the Hall resistivity data suggests additional contributions to the data apart from the OHE and AHE, similar to what was observed in previous reports Ran et al. (2023); Ni et al. (2021); Zhang et al. (2024a). Similar peaks in the field-dependent Hall resistivity of the nanoflake MST device (Device D1 of SM 32) are observed in Fig. 1(d). If the observed low-field nonlinearity were dominated by field-induced changes in carrier density or mobility, one would expect a comparable nonlinearity to persist or evolve continuously into the high-field regime. However, it is clear from Fig. 1(c),(d), depicts a strictly linear field dependence at high fields for both bulk and nanoflake devices. This indicates possible contributions from the THE.
@@
To further examine the origin of the sharp peak observed in the Hall resistivity data, we measure the field dependence of at different temperatures for both bulk and nanoflake device of MST, as shown in Fig. 2(a) and 2(b), respectively. Figures 2(a),(b) confirms that the peak in remains confined to the same low-field regime over a broad temperature range. The high-field Hall slope also remains linear and unchanged ruling out the possibilty of carrier-based nonlinear Hall mechanisms which typically strengthen at higher fields which is not the case here. For bulk, the peak in the Hall resistivity data persists up to to 70 K, while for the device, it starts to die out above 30 K. The occurrence of this peak in the curve for bulk MST has been attributed to COC in previous reports Zhang et al. (2024a). This COC is known to induce an orbital moment (MCOC) along the c axis Zhang et al. (2022), that underpins the well-known CMR in bulk MST. This MCOC produces a magnetic field Bc in the real space, which is directed along the c-axis. When B c, it adds up to the external field applied and influences the Hall resistivity Zhang et al. (2024a).
@@
where , and , refers to ordinary, anomalous and topological Hall contribution to , respectively. Ms refers to saturation magnetization, while R0 and are the ordinary and anomalous Hall coefficients, respectively. The ordinary Hall component usually dominates as a high-field linear component of the data and is separated out by calculating R0 using a linear fit at high fields. For bulk, is isolated by fitting with the magnetization curve. For the device, it is separated by fitting with a step-like function similar to Bera and Mukhopadhyay (2025).
Although in materials exhibiting CMR, R0 and can be field dependent, however, for MST, the Hall resistivity at sufficiently high magnetic fields is strictly linear, despite the presence of CMR. This allows a robust experimental determination of an effective R0 in the high-field regime where the non-linear Hall effect is completely absent. The anomalous Hall term depends explicitly on magnetization along B c, which increases monotonically with field and does not saturate even at high fields, evident from the inset of Fig. 1(a). The additional Hall contribution appears only in a narrow low-field window and vanishes well before magnetization saturation, indicating that even if were allowed to vary smoothly with field, the product cannot generate a sharp, symmetric peak localized near zero field.
@@
Figure 2(c) and (d) shows the fitted curves for bulk and device MST, respectively. The red solid line shows the ordinary and anomalous contributions in with B c at 5 K in both cases. The sharp peak region, in excess of the ordinary and anomalous Hall contribution, is attributed to the THE and is extracted by subtracting the curves by the red solid line data. The contour plots of the extracted THE amplitude for both the bulk and nanoflake device are shown in Fig. 2(e),(f), respectively. From the figures, it is evident that the THE amplitude for the bulk persists up to a higher temperature when compared to the nanoflake device. This could be because of the COC state being more vulnerable to thermal fluctuations when one approaches nanoscale thickness limit compared to the bulk. The THE amplitude mimicking the sensitivity of the COC state to sample thickness suggests that the former might be governed by the latter.
@@
To ascertain the causal connection of the COC and the observed THE in MST, we perform the magnetic field dependence of the Hall resistivity as a function of the applied current. Since it is well known that the strength of the COC decreases with increasing current Zhang et al. (2024a), the relation of to the current will be crucial in studying its actual origin. Figure 3 depicts the extracted THE amplitudes as a function of temperature and magnetic field for different applied currents for both bulk [Fig 3(a) - (e)] and device [Fig 3(f) - (j)], respectively. As the measurements of bulk and device were performed at current values with orders of magnitude difference, we employ current densities for meaningful comparison. It is evident from the figure that upon increasing the current, the THE amplitudes decrease for both the bulk and the nanoflake device. This confirms that the THE signal is not an artifact of polaronic effect. Polaronic effects typically strengthen with increased carrier activation or Joule heating, i.e., with increasing current. Here, we see a complete opposite trend confirming the connection between THE and COC.
@@
For bulk MST, the THE amplitude remains non-zero even up to 70 K when the applied current density is small (5.1 A cm-2). As the current density increases to 51 A cm-2, the region of the non-zero THE amplitude seems to split into two regimes of topological Hall signals [Fig. 3(b)], which shrink and eventually vanishe upon increasing the current density to 1020 A cm-2 as seen in Fig. 3(e), similar to that reported in Zhang et al. (2024a). This is expected as the COC is stronger for lower currents (leading to a lower current density) but on increasing the current (or current density) it starts to weaken. At high current density, the COC state () is destroyed completely, making way for a trivial state (). Similar trend is observed to be followed by the THE signal of the bulk MST crystals, indicating that the two might be interconnected.
Although the THE signal from the nanoflake device of MST replicates the current dependence in case of the bulk MST, the device shows the disappearance of the THE behaviour at a much lower current density than the bulk, as is evident from, Fig. 3(f) - (j). This reflects the reduced robustness of the COC state under reduced dimensionality. The lower critical current density for the disappearance of the THE contribution in the nanoflake device originates from the increased vulnerability of the COC state under reduced dimensionality. Thickness reduction limits the coherence volume of orbital currents, enhances surface-induced orbital dephasing, increases sensitivity to local electric fields and thermal fluctuations, and weakens interlayer phase locking. As a result, the orbital-coherent state collapses at lower current density in nanoflake devices compared to bulk crystals, while preserving the same underlying physical mechanism. This systematic reduction of the critical current density with thickness is a hallmark of a fragile, coherence-based orbital state rather than a conventional transport effect.
@@
From previous reports, it is well-established that the CMR portrayed by the MST is governed primarily by the COC Zhang et al. (2022, 2024a); Das and Mukhopadhyay (2025). Thus, drawing a correlation between the THE and the magnetoresistance of MST in bulk and device, respectively, will provide a firm assurance of the COC origin of the THE. Figure 4(a),(b) shows the temperature dependence of MR and relative topological Hall conductivity for the bulk and device of MST, respectively. The MR is calculated as, , where, with being the magnetoresistivity measured at an applied magnetic field of B T, and the zero-field (ZF) resistivity. The hall conductivity is given by , where and are the longitudinal and Hall resistivity, respectively. Similarly, the topological Hall conductivity is given by , where is the extracted topological Hall resistivity.
@@
Thus, we observe that the THE in both bulk and nanoscale device of MST has a very strong correlation with the phenomena of CMR. As the latter is directly influenced by the COC in MST, the former could possibly have the same origin. It has already been reported that the COC develop a moment MCOC which induces a magnetic field Bc in real space, directed along the c-axis. This gives rise to a non-zero local scalar spin chirality, and the moment MCOC acts as the topological orbital moment (TOM). This couples with the local spins of Mn giving rise to spin-chiral interaction (SCI), mediated by SOC. Also, each individual COC can interact with each other, giving rise to a chiral-chiral interaction (CCI) Grytsiuk et al. (2020). The formation of these two topological chiral interactions give rise to a finite spin chirality. Now, the THE contribution to the resistivity comes from two terms: , where refers to the transverse charge current due to chiral orbital textures and comes from chiral spin textures Denisov et al. (2018). Hence, systems with strong dominating chiral orbital textures, like COC, can also show THE, which might be the reason behind the observed THE in MST.